# A Survey on the Applications of Zero-Knowledge Proofs

### v1

@@ III-D2 Methodology
Both MSM and NTT can be accelerated on GPUs, particularly MSM through an algorithm called ‘Pippenger’. This process involves rewriting the computationally intensive tasks from the CPU to the GPU using CUDA or OpenCL, allowing the code to be compiled and executed directly on the GPU. For finer-grained acceleration, developers can optimize memory usage by maximizing the use of fast memory and minimizing slow-access memory to reduce costly data transfers, especially between the CPU and GPU. Additionally, optimizing execution configuration by balancing work across multiprocessors, building concurrent kernels, and allocating resources judiciously can maximize hardware utilization. The goal is to parallelize the entire workflow, minimizing sequential execution where different parts depend on each other’s results. Open-source implementations allow developers to quickly start their modifications [59].
FPGAs, or field-programmable gate arrays, offer programmable hardware fabric that can be reconfigured multiple times, cutting manufacturing costs compared to ASICs and providing greater flexibility in hardware resource usage than GPUs. Although optimizing NTTs on GPUs is achievable, frequent data shuffling can lead to significant communication overhead between the GPU and CPU. By implementing the logic directly into the circuit design, FPGAs can potentially perform the task faster. Most open-source implementations for zero-knowledge proofs are written in Rust due to its memory safety and cross-platform compatibility. However, FPGA development tools are typically adapted to C/C++, requiring teams to translate these implementations [60].
GPUs offer fast development times with well-documented frameworks like CUDA and OpenCL and are readily available and cost-effective. However, GPUs are power-hungry, even when exploiting data and thread-level parallelism. In contrast, FPGAs have a more complex development cycle, requiring specialized engineers but allowing for low-level optimizations and providing lower latency, especially for large data streams. FPGAs are more expensive and less readily available than GPUs [58].
ASICs, or application-specific integrated circuits, are customized for particular uses and are permanently etched into silicon, making the design and manufacturing process much more complex and time-consuming compared to FPGAs. Despite this, advancements such as Leo’s new integrated chip for accelerated proof generation demonstrate ongoing developments in this area. ASICs are believed to be the most promising hardware acceleration; however, they still have major barriers to entry. Programmability and logic modifications are difficult on ASICS as they possess write-once business logic, necessitating a complete rebuild of the system for any modifications. Conversely, FPGAs and GPUs can be reprogrammed multiple times, allowing the same hardware to be used across various projects with different proof systems or updates. This reprogrammability makes FPGAs a more versatile alternative compared to ASICs. Additionally, the time required for ASIC design, production, and deployment usually spans 12 to 18 months or more [58].
III-D3 Applications
Ingonyama [61] is a hardware acceleration company that integrates chip design with mathematics and advanced algorithms to enhance the performance of compute-intensive cryptography. They maintain a library called ICICLE, a cryptography library for ZKPs using GPUs. ICICLE implements various cryptographic primitives such as elliptic curve (EC) operations, multi-scalar multiplication (MSM), number theoretic transform (NTT), and the Poseidon hash on GPUs. The Polynomial API offers a framework for polynomial operations for developer convenience. Additionally, ICICLE has bindings for Rust and Golang and integrates with projects like Gnark and EZKL. It can also be run in Google Colab. Ingonyama is advancing for a zero-knowledge processing unit (ZPU) defined as “a versatile and programmable hardware accelerator, designed to address the emerging needs of ZKP processing.”
Cysic [62] is a ZK accelerator focused on developing ASICs and their accelerated zkVM. The system architecture features an executor responsible for executing programs, hardware for controlling and distributing segments, and a configurable number of specialized chips to generate ZK proofs for each segment program. Leveraging its expertise in ASIC design and GPU engineering, Cysic aims to overcome challenges by offering ZK compute-as-a-service to various ZK projects. They aim to produce ASICs that specifically target MSMs, NTTs, and other general operations, while maintaining flexibility to adapt to many proving systems. Cysic’s ongoing efforts concentrate on specialized ASIC design for real-time ZK proof generation.
Fabric Cryptography [63] introduces The Fabric Verifiable Processing Unit (VPU), a processor designed for cryptography applications ranging from ZKP to FHE. The VPU features a custom instruction set architecture tailored for next-generation cryptography, including ZKP, FHE, MPC, and other algorithms. It offers acceleration for MSM, NTT, polynomial evaluation, as well as Poseidon (1, 2), Blake, and other hash functions. The VPU supports multiprecision vector lanes up to 384-bit and includes a RISC-V core for enhanced programmability. It supports PCIe 4.0 x16 lanes, providing up to 256 Gbps per chip, and is equipped with high-bandwidth DRAM. In addition, Fabric Cryptography offers a PCIe card featuring 3x FC1000 chips for parallel ZK proof generation, a PCIe interface for comprehensive on-chip proving and encryption, and DRAM for recursive ZK proof generation and mining workloads. They also provide server systems and data centers to support larger workloads.
Irreducible [64] offers proving as a service designed for scalability, powered by FPGA-accelerated server clusters. Irreducible supports popular proof systems such as Plonky2 and Polygon zkEVM, with plans to support next-generation systems like Binius and Plonky3. Their FPGA-accelerated server clusters are specifically designed for cryptographic computation at scale. By connecting their FPGAs directly using AMD’s high-speed Aurora protocol, they minimize unnecessary data transfers between the CPU and FPGAs. Operating independently of public cloud infrastructure providers like Amazon Web Services and Google Cloud Platform provides redundancy essential for the decentralized networks they support. Irreducible also features a fully-pipelined FPGA architecture for ZKP-friendly Merkle trees using the Poseidon hash and NTT.
Supranational [65] offers hardware-accelerated cryptography for verifiable and confidential computing. They provide BLST, an IETF-compliant BLS12-381 signature library focused on security and performance. They also implement a simple API for generating VDF and SNARK proofs, powered by fast, open-source implementations running on a high-availability cloud. Additionally, Supranational is developing Sppark, an arbitrary-precision arithmetic accelerator for cryptographic operations, including VDFs, SNARKs, polynomial commitments, and accumulators.
IV Blockchain Applications
@@
Supply Chain/Enterprise Blockchain QEDIT, zk-BeSC
@@
TABLE V: Overview of blockchain-based applications of Zero-Knowledge Proofs
IV-A Layer 1 Blockchains
(a) Layer-1 Data Compression
(b) Layer-1 Privacy
@@ IV-A Layer 1 Blockchains
Figure 8: Comparison of Layer-1 Data Compression and Privacy Approaches using ZKPs
IV-A1 Motivation and Definition
@@
Layer 1 blockchains, such as Ethereum and Bitcoin, are foundational blockchain networks that provide a transparent and immutable ledger for data storage. However, these platforms face challenges in privacy and scalability. The emergence of zero-knowledge proofs (ZKPs) has induced the development of novel Layer 1 solutions that prioritize privacy and data efficiency. These blockchains integrate ZKPs directly into their base layers, offering a dual advantage: enhanced privacy through transaction concealment and improved scalability via data compression, as illustrated in Figure 8. The primary motivation for ZKP-based Layer 1 blockchain is to bridge the gap between the inherent transparency of conventional blockchains with the increasing demand for data privacy and efficient on-chain data management.
IV-A2 Methodology
@@
Layer-1 blockchains utilize ZKPs to obscure critical transaction details, such as the identities of parties involved and the transaction amounts. This concealment is achieved through advanced cryptographic techniques, including zk-SNARKs or zk-STARKs, that validate transactions without revealing their underlying data. To address scalability challenges, ZKP-based Layer-1 blockchains employ state compression. This technique utilizes ZKPs to create compact proofs that validate large sets of transactions or state transitions, thereby reducing the data volume required on-chain. In these networks, participating nodes are responsible for generating and verifying ZKPs. This ensures transaction integrity while maintaining privacy. The consensus mechanisms of these blockchains are uniquely designed to incorporate ZKP validation. This integration ensures that only transactions authenticated through ZKPs are confirmed and appended to the blockchain.
IV-A3 Applications
Cryptocurrency Cryptographic Method Purpose
ZCash zkSNARKs Enables the encryption of transaction data to provide privacy by allowing transaction verification without revealing sender, receiver, or amount.
Monero Ring Signatures, Stealth Addresses, Ring Confidential Transactions (RingCT) Ring signatures obscure the sender’s identity, stealth addresses hide the receiver’s address, and RingCT conceals the transaction amount, ensuring privacy for all transactions.
Aleo zkSNARKs Used for creating private smart contracts which allow for verifiable computations without revealing underlying data, enabling privacy in decentralized applications.
TABLE VI: Overview of Cryptographic Methods in Privacy-Centric Cryptocurrencies
@@ IV-A3 Applications
As a pioneering Layer 1 blockchain, ZCash [66] utilizes zk-SNARKs to offer private transactions, where the details of the sender, receiver, and transaction amounts are encrypted. In ZCash, zk-SNARKs enable the encryption of transaction data, ensuring the anonymity of both the sender and receiver, as well as the confidentiality of the transaction amount. This mechanism allows the network to validate transactions without disclosing the underlying data, thereby preserving the privacy of users. ZCash’s implementation of zk-SNARKs is notable for its efficiency, allowing for the verification of transactions in a matter of milliseconds, which is a significant advancement over other cryptographic techniques that were more computationally intensive. The ZCash blockchain leverages a novel cryptographic method known as a “shielded transaction,” where the transaction metadata is encrypted, and zk-SNARKs are used to prove that the transaction does not violate the network’s consensus rules. This approach enables a high degree of privacy while maintaining the integrity and security of the blockchain. ZCash offers optional privacy features allowing users to choose between shielded and transparent transactions, unlike Aleo and Monero, where privacy is mandatory for all transactions as depicted in Table VI.
Aleo [30] is a distinctive Layer 1 blockchain platform that employs zero-knowledge proofs to enhance privacy and scalability in decentralized applications (dApps). It is designed to facilitate the development and deployment of private applications on the blockchain. Aleo achieves this with the help of a unique framework for constructing dApps that can perform computations in a private and verifiable manner. Its adoption of zk-SNARKs based systems allows developers to create applications where users can interact and transact without revealing sensitive information. This system empowers users to maintain control over their data, ensuring privacy and security in their interactions on the blockchain. Aleo’s approach to blockchain architecture is focused on providing a scalable solution that addresses the common limitations associated with public blockchains, such as privacy concerns and throughput bottlenecks.
@@
The Mina Protocol [54] is an innovative Layer 1 blockchain that introduces a unique approach to scalability and privacy through the use of recursive zk-SNARKs. This innovative protocol is designed to maintain the succinctness of a blockchain by keeping the network size constant ( 22kB) regardless of the total number of transactions processed. This is achieved through the recursive composition of zk-SNARKs, which allows each new block to contain proof of the validity of the entire blockchain history. As a result, the Mina Protocol maintains a drastically reduced blockchain size compared to traditional blockchains, enhancing scalability and usability.
IV-B Layer 2 Scaling
IV-B1 Motivation and Definition
@@
As aforementioned, ZKP can be used for the property of data succinctness, especially when it comes to blockchain validity proofs. A significant pain point of blockchains is the scarcity and cost of block space, which can be solved using Layer 2 scaling solutions such as rollups. The most prominent types of rollups are optimistic rollups and ZK rollups, but the latter will be our main focus. In ZK rollups, zero-knowledge proofs are used to succinctly prove the validity of state changes to a Layer 1 blockchain without requiring validator nodes on a Layer 1 chain to execute those transactions. In essence, the Layer 2 rollup becomes a cheap modular execution layer that benefits from the security and decentralization of Layer 1, and this enables blockchains to scale by significantly reducing usage costs. Some ZK rollups can also provide privacy-preserving properties using zero-knowledge proofs.
IV-B2 Methodology
@@
In Layer 2 rollups [67] (all of which achieve finality on Ethereum), computation is handled outside of Layer 1, and only state changes, deposits, and withdrawals are posted to Layer 1 through the rollup smart contract. This Layer 1 smart contract contains and maintains a state root, which is the Merkle tree root of the state of the rollup, including accounts and balances. A rollup sequencer is an entity, which could range from a single server to a decentralized network of nodes, which orders transactions, produces L2 blocks, and adds rollup transactions to the ZK-rollup contract with a ZK validity proof. The sequencer publishes a batch of highly compressed transactions as well as the previous state root and the new state root after processing all transactions in the batch. Using the ZK validity proof provided by the sequencer, the rollup contract checks that the new state root is valid, then swaps the old root for the new one. The transaction batches published by sequencers are written to Ethereum in the form of encoded function calls, stored either in the calldata of the EVM, which is a data area used to pass arguments to a function and does not modify the blockchain’s state, or in temporary data storage locations called blobs (post-EIP-4844). This serves as a cheap way to store data on-chain, making it possible for individuals to re-construct the state of the rollup using such compressed transactions. Since the Layer 1 rollup contract can quickly verify a zk-SNARK or zk-STARK proof on any amount of large computation, computation is almost fully offloaded to the Layer 2. Through these methodologies, ZK rollups minimize the space and computation restraints of Layer 1 Ethereum, which simply validates state changes and provides inherent decentralization and security [67].
Figure 9: Generalized ZK Rollup Architecture, redrawn based on [68]
@@
Because Layer 2 scaling solutions are built on top of smart contract-based blockchains, specifically Ethereum, rollups need to provide proofs for state transitions on the Ethereum Virtual Machine (EVM), which is Ethereum’s Turing-complete computation engine. Since ZK-rollups must attest to the correctness of computations with ZKP before posting batches to Layer 1, they must go through extra steps in code execution which the EVM does not do. Rather than loading smart contract bytecode into the execution layer and simply posting the result of those computations, ZK-rollups must generate validity proofs for each transaction’s state transition.
Because EVM opcodes are designed for general-purpose computations, proving EVM computations in ZK circuits is too resource intensive and complex. Consequently, it is very difficult to ensure EVM-compatibility in ZK rollups, resulting in varying architectural designs:
@@
  • zkEVM: A zkEVM is a solution which embeds zero-knowledge proofs into EVM smart contract execution by recreating existing EVM opcodes for proving in circuits. A zkEVM computes state transitions just like the typical EVM, however it creates ZK validity proofs to verify the correctness at every operation, including state changes and computations.
@@
  • Custom VMs: Some approaches involve creating new high-level or intermediate languages and virtual machines that are more amenable to ZK proofs but can still support EVM-like operations.
@@
zkEVMs allow easier execution of Solidity smart contracts, while custom VMs may require developers to write smart contracts in some other smart contract language or modify their existing EVM-based implementations due to ZK limitations.
IV-B3 Applications

### v2

@@ 3.4.2 Methodology
Both MSM and NTT can be accelerated on GPUs, particularly MSM using the Pippenger algorithm. This process involves rewriting the computationally intensive tasks from the CPU to the GPU using CUDA or OpenCL, allowing the code to be compiled and executed directly on the GPU. For finer-grained acceleration, developers can optimize memory usage by maximizing fast memory and minimizing slow-access memory to reduce costly data transfers, especially between the CPU and GPU. Additionally, optimizing execution configuration by balancing work across multiprocessors, building concurrent kernels, and allocating resources judiciously can maximize hardware utilization. The goal is to parallelize the entire workflow, minimizing sequential execution where different parts depend on each other’s results. Open-source implementations allow developers to quickly start their modifications [49]. FPGAs, or field-programmable gate arrays, offer programmable hardware fabric that can be reconfigured multiple times, cutting manufacturing costs compared to ASICs and providing greater flexibility in hardware resource usage than GPUs. Although optimizing NTTs on GPUs is achievable, frequent data shuffling can lead to significant communication overhead between the GPU and CPU. By implementing the logic directly into the circuit design, FPGAs can potentially perform the task faster. Most open-source implementations for zero-knowledge proofs are written in Rust due to its memory safety and cross-platform compatibility. However, FPGA development tools are typically adapted to C/C++, requiring teams to translate these implementations [90]. Table 6 provides a compact CPU/GPU/FPGA/ASIC comparison across the dominant proving kernels. Acceleration changes which applications are feasible. High-throughput rollups, zkEVM-style proving, and large ZKML inference proofs often become practical only when MSM/NTT/hashing kernels are accelerated. However, even with acceleration, systems can remain bottlenecked by memory footprint, proof size, and data-availability costs in blockchain deployments.
Table 6: Hardware acceleration comparison for common ZK kernels.
Kernel / constraint CPU GPU FPGA ASIC
MSM baseline, flexible strong parallelism low-latency pipelines highest throughput potential
FFT/NTT optimized libraries good if bandwidth fits strong if streaming fits strong if specialized
Hashing / Merkle baseline strong for batched hashing very strong for pipelines very strong for fixed hash
Memory bandwidth cache-limited HBM helps, PCIe transfers hurt on-chip bandwidth helps design-specific
3.4.3 Projects
Ingonyama [45] is a hardware acceleration company that integrates chip design with mathematics and advanced algorithms to enhance the performance of compute-intensive cryptography. They maintain a library called ICICLE, a cryptography library for ZKPs using GPUs. ICICLE implements various cryptographic primitives such as elliptic curve (EC) operations, multi-scalar multiplication (MSM), number theoretic transform (NTT), and the Poseidon hash on GPUs. Cysic [22] is a ZK accelerator focused on developing ASICs and their accelerated zkVM. The system architecture features an executor responsible for executing programs, hardware for controlling and distributing segments, and a configurable number of specialized chips to generate ZK proofs for each segment program. Fabric Cryptography [28] introduced The Fabric Verifiable Processing Unit (VPU), a processor designed for cryptography applications ranging from ZKP to FHE. The VPU features a custom instruction set architecture tailored for next-generation cryptography, including ZKP, FHE, MPC, and other algorithms. It offers acceleration for MSM, NTT, polynomial evaluation, as well as Poseidon, Blake, and other hash functions. The VPU supports multiprecision vector lanes up to 384-bit and includes a RISC-V core for enhanced programmability.
Irreducible [46] offers proof-as-a-service, designed for scalability and powered by FPGA-accelerated server clusters. Irreducible supports popular proof systems such as Plonky2 and Polygon zkEVM, with plans to support next-generation systems like Binius and Plonky3. Their FPGA-accelerated server clusters are specifically designed for cryptographic computation at scale. Supranational [101] offers hardware-accelerated cryptography for verifiable and confidential computing. They provide BLST, an IETF-compliant BLS12-381 signature library focused on security and performance.
4 Blockchain Applications
Table 7: Overview of blockchain-based applications of Zero-Knowledge Proofs
@@ 4 Blockchain Applications
Supply Chain/Enterprise Blockchain QEDIT, zk-BeSC, Sahai et al.
@@
4.1 Layer 1 Blockchains
@@
Figure 6: Comparison of Layer-1 Data Compression and Transaction Privacy Approaches using ZKPs
4.1.1 Motivation and Definition
@@
Layer 1 blockchains, such as Ethereum and Bitcoin, are foundational blockchain networks that provide a transparent and immutable ledger for data storage. However, these platforms face challenges in privacy and scalability. The emergence of zero-knowledge proofs (ZKPs) has induced the development of novel Layer 1 solutions that prioritize privacy and data efficiency, as depicted in Figure 6. These blockchains integrate ZKPs directly into their base layers, offering a dual advantage: enhanced privacy through transaction concealment and improved scalability via data compression. In practice, privacy-first Layer 1s select a proving system based not only on proof size and prover cost, but also on the setup model (trusted vs. transparent) and whether recursive composition is a core requirement for keeping verification lightweight over long histories. The primary motivation for ZKP-based Layer 1 blockchains is to reconcile the inherent transparency of conventional blockchains with the increasing demand for data privacy and efficient on-chain data management.
4.1.2 Methodology
@@
Layer-1 blockchains utilize ZKPs to obscure critical transaction details, such as the identities of parties involved and the transaction amounts. This concealment is achieved through advanced cryptographic techniques, including zkSNARKs or zkSTARKs, that validate transactions without revealing their underlying data. To address scalability challenges, ZKP-based Layer-1 blockchains employ state compression. This technique uses ZKPs to produce compact proofs that verify large sets of transactions or state transitions, thereby reducing the on-chain data volume required. In these networks, participating nodes are responsible for generating and verifying ZKPs. This ensures transaction integrity while maintaining privacy. The consensus mechanisms of these blockchains are uniquely designed to incorporate ZKP validation. This integration ensures that only transactions authenticated through ZKPs are confirmed and appended to the blockchain. However, these benefits introduce operational complexity, such as user UX can degrade due to wallet and key-handling requirements, proving latency can impact transaction confirmation time, especially on constrained devices, and key management becomes a first-class security concern, e.g., circuit upgrades, verification keys, and setup artifacts when applicable.
4.1.3 Applications
@@
ZCash [107], one of the earliest privacy-focused Layer 1 blockchains, uses zkSNARKS to enable private transactions, where the identities of the sender and receiver, and the transaction amounts, are encrypted. In ZCash, zkSNARKS enable the encryption of transaction data, ensuring the anonymity of both the sender and receiver, as well as the confidentiality of the transaction amount. This mechanism allows the network to validate transactions without revealing the underlying data, thereby preserving users’ privacy. ZCash’s implementation of zkSNARKS is notable for its efficiency, allowing for the verification of transactions in a matter of milliseconds, which is a significant advancement over other cryptographic techniques that were more computationally intensive. ZCash uses shielded transactions, in which transaction metadata is encrypted and zkSNARKS are used to prove compliance with consensus rules without revealing the underlying data. This approach enables a high degree of privacy while maintaining the integrity and security of the blockchain. ZCash offers optional privacy features allowing users to choose between shielded and transparent transactions, unlike Aleo, where privacy is mandatory for all transactions.
@@
Aleo [4] is a distinctive Layer 1 blockchain platform that employs zero-knowledge proofs to enhance privacy and scalability in decentralized applications (dApps). It achieves this through a zkSNARK framework for constructing dApps that can perform computations in a private and verifiable manner. The Mina Protocol [60] is a Layer 1 blockchain that addresses scalability and privacy through recursive zkSNARKS. This protocol maintains a very small constant-size proof of chain validity through recursive composition, regardless of the total number of transactions processed. This is achieved through the recursive composition of zkSNARKS, which allows each new block to contain proof of the validity of the entire blockchain history. As a result, the Mina Protocol maintains a drastically reduced blockchain size compared to traditional blockchains, enhancing scalability and usability.
4.2 Layer 2 Scaling
4.2.1 Motivation and Definition
@@
As mentioned above, ZKP can be used for data succinctness, especially for blockchain validity proofs. A significant pain point for blockchains is the scarcity and cost of block space, which can be addressed through Layer 2 scaling solutions such as rollups. The most prominent types of rollups are optimistic rollups and ZK rollups, but we will focus on the latter. More precisely, validity rollups use succinct proofs to attest to correct state transitions: zero-knowledge privacy is optional and protocol-specific. In ZK rollups, zero-knowledge proofs are used to succinctly prove the validity of state changes to a Layer 1 blockchain without requiring validator nodes on a Layer 1 chain to execute those transactions. In essence, the Layer 2 rollup becomes a cheap, modular execution layer that benefits from the security and decentralization of Layer 1, enabling blockchains to scale by significantly reducing operational costs.
4.2.2 Methodology
@@
In Layer 2 rollups [27], all of which achieve finality on Ethereum Layer 1, heavy computation is handled outside the chain, and only state changes, deposits, and withdrawals are posted to Layer 1 through the rollup smart contract. This Layer 1 smart contract maintains a state root, the Merkle tree root of the rollup’s state, including accounts and balances. A rollup sequencer is an entity that could range from a single server to a decentralized network of nodes that orders transactions, produces L2 blocks, and adds rollup transactions to the ZK-rollup contract with a ZK validity proof. The sequencer publishes a batch of highly compressed transactions, along with the previous state root and the new state root, after processing all transactions in the batch. Using the ZK validity proof provided by the sequencer, the rollup contract checks that the new state root is valid, then swaps the old root for the new one. The transaction batches published by sequencers are written to Ethereum in the form of encoded function calls, stored either in the calldata of the EVM, which is a data area used to pass arguments to a function and does not modify the blockchain’s state, or in temporary data storage locations called blobs (post-EIP-4844). This provides a cheap way to store data on-chain, enabling individuals to reconstruct the rollup’s state from these compressed transactions. Since the Layer 1 rollup contract can quickly verify a zk-SNARK or zk-STARK proof on any amount of large computation, computation is almost fully offloaded to the Layer 2. Through these methodologies, ZK rollups minimize the space and computation restraints of Layer 1 Ethereum, which simply validates state changes and provides inherent decentralization and security [27]. In practice, user fees are often dominated by (i) on-chain verification cost (proof verification and any aggregation), and (ii) data availability cost, where calldata and blob-based DA trade off pricing and throughput. Consequently, DA policy (rollup vs. validium vs. volition) and compression strategy can matter as much as the prover itself.
@@
Since Layer 2 scaling solutions are built atop smart contract-enabled blockchains, most notably Ethereum rollups, they must submit proofs of their state transitions to the Ethereum Virtual Machine (EVM), which serves as Ethereum’s Turing-complete computation engine. Since ZK-rollups must attest to the correctness of computations using ZKPs before posting batches to Layer 1, they must take additional steps during code execution, whereas the EVM does not. Rather than loading smart contract bytecode into the execution layer and simply posting the result of those computations, ZK-rollups must generate validity proofs for each transaction’s state transition. Because EVM opcodes are designed for general-purpose computations, proving EVM computations in ZK circuits is too resource-intensive and complex. Consequently, it is very difficult to ensure EVM-compatibility in ZK rollups, resulting in varying architectural designs:
@@
  • zkEVM: A solution that embeds zero-knowledge proofs into EVM smart contract execution by recreating existing EVM opcodes for proving in circuits. It computes state transitions just like the typical EVM. However, it creates ZK validity proofs to verify the correctness at every operation, including state changes and computations.
@@
  • Custom VMs: Some approaches involve creating new high-level or intermediate languages and virtual machines that are more amenable to ZK proofs while still supporting EVM-like operations.
@@
zkEVMs enable easier execution of Solidity smart contracts, whereas custom VMs may require developers to write smart contracts in another smart contract language or modify their existing EVM-based implementations due to ZK limitations.
Table 8: Comparison of representative Layer 2 validity systems.
L2 Proof family Setup EVM level Language/tooling DA mode Indicative proving throughput
zkSync Era SNARK trusted EVM-compatible Solidity/Vyper + custom compiler rollup/validium medium
Polygon zkEVM SNARK trusted EVM-equivalent Solidity (no custom compiler) rollup/validium medium
Scroll SNARK trusted EVM-compatible Solidity (minimal changes) rollup medium
Linea SNARK trusted EVM-equivalent Solidity (standard tooling) rollup medium
StarkNet STARK transparent custom VM Cairo toolchain rollup/validium/volition high
Aztec SNARK trusted custom VM Noir/ACIR toolchain rollup low–medium
4.2.3 Applications